OpenStudy (anonymous):

The speed of train A is 8 mph slower than the speed of train B. Train A travels 230 miles in the same time it takes train B to travel 270 miles. Find the speed of each train.

OpenStudy (ajprincess):

Let the speed of train A be x and the speed of train B be y. x+8=y \[speed=\frac{distance}{time}\] \[time=\frac{distance}{speed}\] For train A \[time=\frac{distance}{speed}\] \[time=\frac{230}{x}\] For train B \[time=\frac{distance}{speed}\] \[time=\frac{270}{y}\] \[time=\frac{270}{x+8}\] Since the time taken are equal \[\frac{230}{x}=\frac{270}{x+8}\] Solve for x to find the speed of train A. Substitute the value of x in x+8=y to find the speed of train B.